OpenStudy (sorryimfrankie):

Expand the following using either the Binomial Theorem or Pascal’s Triangle.

OpenStudy (sorryimfrankie):

Expand the following using either the Binomial Theorem or Pascal’s Triangle. (2x + 4)^3 Can someone walk me through this?

OpenStudy (dayakar):

(2x)^3+3.(2x)^2.4+3.(2x).4^2+4^3

OpenStudy (dayakar):

=8x^3+3.4.4.x^2+3.2.16.x+64 =8x^3+46x^2+96x+64

OpenStudy (dayakar):

is it clear

OpenStudy (sorryimfrankie):

I'm not sure what the periods are in there for?

OpenStudy (dayakar):

it is very simple like (a+b)^3 = a^3+3a^2b+3ab^2+b^3

OpenStudy (dayakar):

here a=2x, b=4

OpenStudy (michele_laino):

hint: here are the coefficient of the expansion: 

OpenStudy (michele_laino):

coefficients*

OpenStudy (sorryimfrankie):

Yes, I have (a + b)^3 = 1a^3b^0 + 3a^2b^1 + 3a^1b^2 + 1a^0b^3 Is that right?

OpenStudy (michele_laino):

yes! It is right!

OpenStudy (sorryimfrankie):

Awesome, now can you tell me if I plugged everything in correctly? (a + b)^3 = 1(2x)^3(4) + 3(2x)^2(4)^1 + 3(2x)^1(4)^2 + 1(2x)(4)^3

OpenStudy (michele_laino):

no, since the correct expression is: \[\Large {\left( {2x} \right)^3} + 3{\left( {2x} \right)^2} \cdot 4 + 3 \cdot \left( {2x} \right) \cdot 4 + {4^3}\]

OpenStudy (sorryimfrankie):

Oh, thank you! That is so much simpler. Are there any other steps I have to take after that?

OpenStudy (michele_laino):

oops.. I have made a typo, here is the right expression: \[\Large {\left( {2x} \right)^3} + 3{\left( {2x} \right)^2} \cdot 4 + 3 \cdot \left( {2x} \right) \cdot {4^2} + {4^3}\]

OpenStudy (dayakar):

@sorryimfrankie in the last term 1.(2x)^0.4^3

OpenStudy (michele_laino):

after that expression, you have to compute the subsequent quantities, using the rules of algebra: \((2x)^3=...?\) \((2x)^2=...?\) \(4^2=...?\) and, \(4^3=...?\)

OpenStudy (sorryimfrankie):

(2x)^3 = (6x) (2x)^2 = (4x) 4^2 = 16 4^3 = 64 Correct?

OpenStudy (michele_laino):

the last 2 expression are correct. for the first two expressions, please note that: \[\huge \begin{gathered} {\left( {2x} \right)^3} = {2^3} \cdot {x^3} = ...? \hfill \\ {\left( {2x} \right)^2} = {2^2} \cdot {x^2} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

expressions*

OpenStudy (dayakar):

(2x)^3= 2x.2x.2x=8x^3 like this do remaining terms i already did above

OpenStudy (sorryimfrankie):

I'm sorry dayakar, you're just confusing me

OpenStudy (michele_laino):

hint: \[\huge {\left( {2x} \right)^3} = {2^3} \cdot {x^3} = 8 \cdot {x^3} = 8{x^3}\]

OpenStudy (dayakar):

ok, whom you are following

OpenStudy (sorryimfrankie):

Sorry Michele, I didn't see your last response. The 4th one would be 4x^2, right?

OpenStudy (sorryimfrankie):

Sorry if I'm not getting all of this, my medication is making me kind of slow at things

OpenStudy (michele_laino):

yes! correct!

OpenStudy (sorryimfrankie):

Oh awesome, thanks so much! And is that all I have to do?

OpenStudy (dayakar):

@Michele_Laino i'm sorry ,plz continue

OpenStudy (michele_laino):

next we can write this: \[\Large \begin{gathered} {\left( {2x} \right)^3} + 3{\left( {2x} \right)^2} \cdot 4 + 3 \cdot \left( {2x} \right) \cdot {4^2} + {4^3} = \hfill \\ \\ = 8{x^3} + 3 \cdot \left( {4{x^2}} \right) \cdot 4 + 3 \cdot \left( {2x} \right) \cdot 16 + 64 \hfill \\ \end{gathered} \] now, what is: \[\Large 3 \cdot \left( {4{x^2}} \right) \cdot 4=...?\] and what is: \[\Large 3 \cdot \left( {2x} \right) \cdot 16=...?\]

OpenStudy (michele_laino):

ok! @dayakar

OpenStudy (michele_laino):

hint: we can write this: \[\Large 3 \cdot \left( {4{x^2}} \right) \cdot 4 = 3 \cdot 4 \cdot 4 \cdot {x^2} = ...?\]

OpenStudy (sorryimfrankie):

48x^2 and 96x?

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

so the final answer is: \[\Large \begin{gathered} {\left( {2x} \right)^3} + 3{\left( {2x} \right)^2} \cdot 4 + 3 \cdot \left( {2x} \right) \cdot {4^2} + {4^3} = \hfill \\ \hfill \\ = 8{x^3} + 3 \cdot \left( {4{x^2}} \right) \cdot 4 + 3 \cdot \left( {2x} \right) \cdot 16 + 64 = \hfill \\ \hfill \\ = 8{x^3} + 48{x^2} + 96x + 64 \hfill \\ \end{gathered} \]

OpenStudy (sorryimfrankie):

Awesome! I guess that wasn't too hard after all. Thank you so much! :)

OpenStudy (michele_laino):

:)